OpenStudy (anonymous):

Find the area Triangle=10 2/3, h=12 5/6 @swaggd2

OpenStudy (anonymous):

A = (1/2) * b * h

OpenStudy (anonymous):

this is for me right?

OpenStudy (anonymous):

Yeah ,

OpenStudy (phi):

according to the formula you multiply 3 numbers \[ \frac{1}{2} \cdot 10 \frac{2}{3} \cdot 12 \frac{5}{6} \] one way to do this is first change the mixed numbers to improper fractions can you do that ?

OpenStudy (anonymous):